# The Cooperative Network Architecture: Learning Structured Networks as Representation of Sensory Patterns

### v3

@@ Abstract
Keywords net fragments  $\cdot$ nets  $\cdot$ neural code  $\cdot$ pattern recognition  $\cdot$ computer vision  $\cdot$ neural networks  $\cdot$ machine learning
@@ 1 Introduction
Existing approaches attempt to address these challenges through methods such as feature or sample engineering (Amirian et al., 2018; Moosavi-Dezfooli et al., 2016), domain adaptation (Csurka, 2017; Sager et al., 2022), and data augmentation (Carlucci et al., 2019; Tuggener et al., 2024). However, these techniques largely operate within the conventional paradigm of representing objects as static patterns of neuronal activations, often requiring large datasets and extensive supervision to achieve generalization.
In contrast, we propose a new paradigm by introducing the Cooperative Network Architecture (CNA), a framework in which object representations emerge from dynamic interactions among recurrently connected neurons. Rather than encoding objects as distinct activation vectors, CNA organizes sensory information into “nets” – structured networks of mutually supporting neurons. These nets are assembled from overlapping “net fragments,” which encode frequently co-occurring feature constellations and are acquired through Hebbian plasticity (Hebb, 1949). This compositional structure supports robust recognition, denoising, and figure completion through cooperative dynamics without external supervision.
@@
Our results show that CNA offers a biologically inspired computational framework that integrates local and global information processing. It provides a foundation for more robust and adaptable perception systems, aligning with theoretical proposals for dynamic, network-based neural coding (von der Malsburg, 2018; von der Malsburg et al., 2022). By establishing the neural code and core computational mechanisms within a single representational area, this work provides the essential groundwork for future architectures in which multiple areas cooperate to support invariant object recognition – thus advancing toward the full potential envisioned in network-based neural coding models of perception.
@@ 2 The Net Fragment Framework
Before discussing the computational framework, we first introduce its building blocks: neurons, activations, nets, and net fragments. These definitions provide the basis for the discussions in the subsequent sections. For simplicity, we limit the definitions to a single area even though it could be extended to multiple inter-connected areas.
@@
A neuron $i$ is denoted by $n_{i}$. Each neuron $n_{i}$ has a local neighborhood, denoted by $N_{r}(n_{i})$, which is the set of neurons $n_{j}$ such that the distance $d(n_{i},n_{j})$ between $n_{i}$ and $n_{j}$ is less than or equal to a threshold radius $r>0$, i.e.,
$N_{r}\left(n_{i}\right)=\{n_{j}\mid d(n_{i},n_{j})\leq r\}.$ (1)
@@
$\forall w_{i,j}\quad n_{j}\in N_{r}(n_{i})$ (2)
Additionally, we limit the weights to a range between 00 and $1$, allowing neuron $n_{i}$ only to positively influence $n_{j}$:
@@
A net fragment is defined as the set of neurons that provide direct activation support to a given neuron. In the context of this work, a neuron $n_{j}$ is said to provide support to another neuron $n_{i}$ if two conditions are met: (1) there exists a high correlation between two neurons $n_{j}$ and $n_{i}$ such that $n_{j}\sim n_{i}\Rightarrow w_{ji}>0$; and (2) neuron $n_{j}$ was active at the previous time step, i.e., $y_{j,t-1}=1$. Since synaptic weights are bounded within the interval $w_{j,i}\in[0,1]$, the degree of support a neuron can provide is likewise restricted to this range.
@@
The decision of whether a neuron becomes active is based on the amount of support it receives from within the net fragment. We call this accumulated input the membrane potential at timestep $t$, which is a function $\mathcal{M}$ of the net fragment $\mathcal{F}_{t-1}(n_{i})$ of a neuron $n_{i}$ at timestep $t-1$.
$a^{\mathcal{M}}_{i,t}=\mathcal{M}\left(\mathcal{F}_{t}(n_{i})\right)$ (5)
The membrane potential $a^{\mathcal{M}}_{i,t}\in\mathbb{R}$ is large when many other neurons provide support, meaning that they “confirm” the plausibility of a neuron’s activity by being active themselves and exhibiting a connection that has been strengthened through previous mutual activity and Hebbian updates.
The activation of a neuron at timestep $t$ is a function of the membrane potential $a_{i,t}$ and environmental conditions $\mathcal{E}$:
$y_{i,t}=\mathcal{A}\left(a_{i,t},\mathcal{E}\right)\in\{0,1\}$ (6)
The environment conditions $\mathcal{E}$ include an attenuation signal, a constraint that selectively suppresses weak excitatory inputs by requiring neurons to accumulate a sufficiently large membrane potential before firing.
@@
Neurons are not only influenced by direct connections within net fragments but also by indirect connections (e.g., if $n_{i}$ supports $n_{j}$ and $n_{j}$ supports $n_{k}$, then $n_{i}$ indirectly supports $n_{k}$). A set of directly and indirectly connected neurons is called a net, implementing the basic structure used in this work to represent objects. For simplicity in the following definitions, we introduce $\mathcal{F}^{*}_{t}(n_{i})$, a net fragment $\mathcal{F}_{t}(n_{i})$ in which $n_{i}$ is active:
@@
To account for both direct and indirect influences on a neuron’s activation, we use a recursive formulation of nets from the perspective of a single neuron. The base case is defined as the immediate net fragment:
@@
which comprises all neurons in the active subgraph that are reachable from $n_{i}$. This recursive formulation illustrates how local net fragments, which can be interpreted as elementary features, are progressively integrated to form higher-order features. Moreover, the explicit time dependence in these definitions underscores the dynamic nature of both the net fragments and the overall network, as they are continuously reassembled in response to the evolving support received from neighboring neurons, and increasing attenuation that prevents neurons with a low membrane potential from firing (see next section).
In the proposed architecture, no symmetry constraint is imposed on the weights, that is, $w_{i,j}\neq w_{j,i}$ in general. However, due to the temporal dynamics of the network, activations tend to stabilize over a longer time horizon $T$, such that for sufficiently large $t$, the activations become approximately stationary, i.e., $y_{i,t}\approx y_{i,t+1}$. Under these conditions, mutual influence between neurons becomes symmetric, meaning that $n_{i}\sim n_{j}$ implies $n_{j}\sim n_{i}$, and almost symmetric weights emerge naturally over time, even in the absence of explicit symmetry constraints.
@@ 3 Cooperative Network Architecture
Figure 2: The dynamics within the proposed CNA: The input image is fed into stage $S1$ to obtain feature activations. The feature activations are then processed by stage $S2$, together with activations from neurons within the same layer.
@@
The CNA comprises two processing stages, depicted as yellow squares in Figure 2. The stage 1 ($S1$) performs fixed (i.e., not learned) feature extraction directly from the image input. The stage 2 ($S2$) implements the core principles of the described computational framework by learning net fragments.
The network dynamics unfold over discrete time steps $t=[0,\dots,T]$. Since the feature extraction stage ($S1$) employs fixed filters, it yields a time-invariant activation pattern for a given input. In contrast, the second stage ($S2$) exhibits dynamic temporal evolution, where internal representations are iteratively refined at each time step ($t\to t+1$). Specifically, we use a progressively increasing global attenuation signal that only allows neurons with a high membrane potential to remain active. During this dynamic processing, a neuron might be deactivated, whereby another neuron’s membrane potential is decreased, triggering a chain reaction that can lead to the deactivation of incoherent nets.
While these refinements occur within the processing cycle of a single sample ($[0,\dots,T]$), Hebbian learning adapts the weights (connections $w_{i,j}$ within net fragments) on longer timescales, wherein synaptic weight modifications occur between different sample-processing episodes, leading to structural plasticity in the network. The following sections discuss these two processing stages.
@@ 3.1 Stage 1: Feature Extraction
The input image is denoted as $\boldsymbol{x}\in\mathbb{R}^{C,H,W}$ (illustrated by the gray box in Figure 2), where $C$ indicates the number of image channels ($1$ for grayscale, $3$ for RGB), while $H,W$ stand for the height and width of the image.
@@
To extract initial activation patterns, we use fixed, non-trainable filters (details see Appendix A). These filters are implemented as standard convolutional kernels, used to obtain the binary activation pattern $\boldsymbol{y}^{(S1)}\in\{0,1\}^{C^{(S1)},H,W}$ as follows:
@@
$y_{i}=\Theta(a_{i})=\begin{cases}1\quad\text{if }a_{i}>0\\ 0\quad\text{otherwise}\end{cases}$ (13)
@@ 3.2 Stage 2: Net Fragments
$S2$ constructs net fragments from the binary feature activations $\boldsymbol{y}^{(S1)}$ of $S1$ and integrates them into coherent, spatially extended nets over time. To ensure stability, single neurons or small groups of neurons are suppressed through an attenuation mechanism, ensuring that only large-scale, consistent activation patterns persist.
@@
Each neuron $n^{(S2)}_{i}$ in stage 2 is connected in two ways: (1) Feedforward connections to the corresponding neuron $n^{(S1)}_{i}$ in stage 1 and to its local neighborhood $Nr(n^{(S1)}_{i})$, and (2) lateral connections to neighboring neurons within stage 2 itself, denoted as:
$Nr\left(n^{(S2)}_{i}\right)=Nr\left(n^{(S1)}_{i}\right)\cup\left\{n_{j}^{(S2)}\mid d(n_{i}^{(S2)},n_{j}^{(S2)})\leq r_{2}\right\}$ (15)
where $r_{2}$ defines the radius of lateral connections. The convolutional architecture realizes these neighborhood structures via convolutional filters, where the filter size directly determines the radius of both feedforward and lateral connections. Using convolutional filters not only implements the connectivity restrictions to local neighborhoods but also allows detection of the same pattern at various positions, enabling learning of net fragments independent of position.
@@
To determine the support a single neuron $n_{i,t}$ receives from the net fragment $\mathcal{F}_{t-1}(n_{i,t-1})$, we first calculate the membrane potential $\mathcal{M}$ by applying convolutional operations on the forward and lateral connections and then taking the sum of these operations:
@@
$\boldsymbol{a}^{\mathcal{M}(S2)}_{c_{\text{out}},t}=\underbrace{\left(\sum_{c_{\text{in}}=0}^{C^{(S1)}}\boldsymbol{W}^{(F)}_{c_{\text{out}},c_{\text{in}}}\star\boldsymbol{y}_{c_{\text{in}}}^{(S1)}\right)}_{\textit{Forward Connection}}+\overbrace{\left(\sum_{c_{\text{in}}=0}^{C^{(S2)}}\boldsymbol{W}^{(L)}_{c_{\text{out}},c_{\text{in}}}\star\boldsymbol{y}_{c_{\text{in}},t-1}^{(S2)}\right)}^{\textit{Recurrent Connection}}.$ (16)

### v4

@@ Abstract
Keywords net fragments  $\cdot$ neural code  $\cdot$ pattern recognition  $\cdot$ computer vision  $\cdot$ representation learning
@@ 1 Introduction
Existing approaches attempt to address these challenges through methods such as feature or sample engineering (Amirian et al., 2018; Moosavi-Dezfooli et al., 2016; Meyer et al., 2025), domain adaptation (Csurka, 2017; Sager et al., 2022), and data augmentation (Carlucci et al., 2019; Tuggener et al., 2024). However, these techniques largely operate within the conventional paradigm of representing objects as static patterns of neuronal activations, often requiring large datasets and extensive supervision to achieve generalization.
In contrast, we propose a new paradigm by introducing the Cooperative Network Architecture (CNA), a model in which object representations emerge from dynamic interactions among recurrently connected neurons. Rather than encoding objects as distinct activation vectors, CNA organizes sensory information into “nets” – structured networks of mutually supporting neurons. These nets are assembled from overlapping “net fragments,” which encode frequently co-occurring feature constellations and are acquired through Hebbian plasticity (Hebb, 1949). This compositional structure supports robust recognition, denoising, and figure completion through cooperative dynamics without external supervision.
@@
CNA offers the conceptual advance of a biologically inspired computational framework that integrates local and global information processing. Our results indicate that it provides a foundation for more robust and adaptable perception systems, aligning with theoretical proposals for dynamic, network-based neural coding (von der Malsburg, 2018; von der Malsburg et al., 2022). In this work, we establish the neural code and core computational mechanisms within a single representational area (a layer of neurons in which nets and their constituent fragments are formed, see below). Such a single area learns input-driven net fragments that bind together co-occurring sensory features, enabling the assembly of coherent perceptual nets. Future architectures will extend this principle to multiple interacting areas, in which additional areas can maintain object-associated net fragments that act as top-down hypotheses. Through reciprocal interactions between input-driven and object-memory areas, such systems would support invariant object recognition and more complex perceptual inference, realizing the broader vision of network-based neural coding models of perception (see our discussion in Section 6).
The remainder of this paper is structured as follows. In Section 2, we present the theoretical foundations of our approach and provide the first formal mathematical definition of nets and net fragments, concepts that have previously been described only in prose (von der Malsburg, 2018; von der Malsburg et al., 2022). Section 3 introduces a concrete computational implementation of this framework. In Section 4, we report experimental results demonstrating the model’s ability to generalize to out-of-distribution inputs and to perform noise suppression. Section 5 situates our approach within the broader literature, contrasting it with related models such as associative memory networks, Boltzmann machines, and modern artificial neural networks. Finally, Section 6 discusses the implications of our findings and outlines how the proposed principles could be extended toward multi-area architectures for object recognition.
@@ 2 The Net Fragment Framework
Before presenting the computational framework, we introduce its foundational components: neurons, activations, nets, net fragments, and competition. These theoretical definitions serve as the basis for the subsequent sections. Implementation details for the introduced formalisms are provided in the next section.
@@
A neuron $i$ is denoted by $n_{i}$. Each neuron $n_{i}$ has a local neighborhood, denoted by the closed ball $\mathcal{B}_{r}(n_{i})$, which is the set of neurons $n_{j}$ such that the distance $d(n_{i},n_{j})$ between $n_{i}$ and $n_{j}$ is less than or equal to a threshold radius $r>0$, i.e.,
$\mathcal{B}_{r}\left(n_{i}\right)=\left\{n_{j}\mid d(n_{i},n_{j})\leq r\right\}.$ (1)
@@
$\forall w_{i,j}\quad n_{j}\in\mathcal{B}_{r}(n_{i})$ (2)
Additionally, we limit the weights to a range between 0 and 111, allowing neuron $n_{i}$ only to positively influence $n_{j}$:
@@
A net fragment is defined as the set of neurons that provide direct activation support to a given neuron. A neuron $n_{j}$ is said to provide support to another neuron $n_{i}$ if two conditions are met: (1) there exists a high correlation between two neurons $n_{j}$ and $n_{i}$ such that $n_{j}\sim n_{i}\Rightarrow w_{ji}>0$; and (2) neuron $n_{j}$ was active at the previous time step, i.e., $y_{j,t-1}=1$. Since synaptic weights are bounded within the interval $w_{j,i}\in[0,1]$, the degree of support a neuron can provide is likewise restricted to this range.
@@
The decision of whether a neuron becomes active is based on the amount of support it receives from within the net fragment. We call this accumulated input the pre-synaptic activity at timestep $t$, which is a function $\mathcal{I}$ of the net fragment $\mathcal{F}_{t-1}(n_{i})$ of a neuron $n_{i}$ at timestep $t-1$.
$\boldsymbol{a}^{\mathcal{I}}_{i,t}=\mathcal{I}\left(\mathcal{F}_{t}(n_{i})\right)$ (5)
The pre-synaptic activity $a^{\mathcal{I}}_{i,t}\in\mathbb{R}$ is large when many other neurons provide support. Since supporting neurons exhibit connections that have been selectively strengthened through previous mutual activity and Hebbian updates within specific contexts, their activity effectively serves to confirm the contextual plausibility of a given neuron’s activation. The very general formulation of $\mathcal{I}$ shows that the pre-synaptic activity can be calculated with any desired function. In this work, we specifically define $\mathcal{I}$ as the weighted sum of prior activations from supporting neurons:
$\boldsymbol{a}^{\mathcal{I}}_{i,t}=\sum_{n_{j}\in\mathcal{F}_{t}(n_{i})}\boldsymbol{w}_{j,i}\cdot\boldsymbol{y}_{j,t-1}.$ (6)
The activation of a neuron at timestep $t$ is determined by the activation function $\mathcal{A}$ that depends on the pre-synaptic activity $a_{i,t}$ and environmental conditions $\mathcal{E}$:
$\boldsymbol{y}_{i,t}=\mathcal{A}\left(\boldsymbol{a}_{i,t},\mathcal{E}\right)\in\{0,1\}$ (7)
The environmental conditions $\mathcal{E}$ include an attenuation signal, a constraint that selectively suppresses weak excitatory inputs by requiring neurons to accumulate a sufficiently large pre-synaptic activity before firing. Again, we here provide a very general formulation, using $\mathcal{A}$ and $\mathcal{E}$, to emphasize that any activation function or environmental mechanism can be used. In this work, we instantiate $\mathcal{E}$ using a combination of normalization (see Appendix B) and attenuation (see Eq. 28), and define $\mathcal{A}$ as the heaviside step function (see Eq. 29).
@@
Neurons are not only influenced by direct connections within net fragments but also by indirect connections (e.g., if $n_{i}$ supports $n_{j}$ and $n_{j}$ supports $n_{k}$, then $n_{i}$ indirectly supports $n_{k}$). A set of directly and indirectly connected neurons is called a net, implementing the basic structure used in this work to represent objects. For simplicity in the following definitions, we introduce $\mathcal{F}^{*}_{t}(n_{i})$, a net fragment $\mathcal{F}_{t}(n_{i})$ in which $n_{i}$ is active as well:
@@
To account for both direct and indirect influences on a neuron’s activation, we use a recursive formulation of nets from the perspective of a single neuron. In the base case, a net $\mathcal{N}_{t}^{1}[n_{i}]$ is defined as the immediate net fragment:
@@
which comprises all neurons in the active subgraph that are reachable from $n_{i}$. This recursive formulation illustrates how support relationships propagate through time: local net fragments $\mathcal{F}_{t}(n_{i})$, which reflect immediate sources of support for a neuron, are combined recursively to form increasingly larger subnetworks $\mathcal{N}_{t}^{r}[n_{i}]$. These progressively include neurons with indirect but recurrent influence on $n_{i}$’s activation. This process reflects how localized evidence (elementary activations) can aggregate into a coherent larger pattern (composite activation), corresponding to a higher-order feature representation in the network. Moreover, the explicit time dependence in these definitions underscores the dynamic nature of both the net fragments and the overall net, as they are continuously reassembled in response to the evolving support received from neighboring neurons, and increasing attenuation that prevents neurons with a low pre-synaptic activity from firing (see next section).
In the proposed architecture, no symmetry constraint is imposed on the weights, that is, $w_{i,j}\neq w_{j,i}$ in general. However, due to the temporal dynamics of the network, activations tend to stabilize over a longer time horizon $T$, such that for sufficiently large $t$, the activations become approximately stationary, i.e., $y_{i,t}\approx y_{i,t+1}$. Under these conditions, the mutual influence between neurons becomes symmetric, meaning that $n_{i}\sim n_{j}$ implies $n_{j}\sim n_{i}$, and almost symmetric weights emerge naturally over time, even in the absence of explicit symmetry constraints.
@@
Competitive Neurons.
One issue with the current formulation is that the mapping
$\mathcal{A}\circ\mathcal{I}:2^{\mathcal{N}}\;\longrightarrow\;\{0,1\}$ (13)
treats any two net fragments that yield the same activation identically. In particular, we say
$\mathcal{F}_{t_{k}}(n_{i})\simeq\mathcal{F}_{t_{l}}(n_{i})\quad\Longleftrightarrow\quad\mathcal{A}\bigl{(}\mathcal{I}(\mathcal{F}_{t_{k}}(n_{i})),\mathcal{E}\bigr{)}=\mathcal{A}\bigl{(}\mathcal{I}(\mathcal{F}_{t_{l}}(n_{i})),\mathcal{E}\bigr{)},$ (14)
meaning different contexts collapse to the same output for $n_{i}$. To break this symmetry and make each neuron context-aware, we introduce competitive neurons.
For each neuron $n_{i}$, We introduce a set of alternatives $\{n_{i}^{1},\dots,n_{i}^{\kappa}\}$, all located at the same spatial position $\left(\mathcal{B}_{r}(n_{i})=\mathcal{B}_{r}(n_{i}^{k})\right)$. These alternatives undergo competition and one of them is permitted to be active within a given context. This exclusivity encourages these competitive neurons to specialize, over time, into representations of different contextual variations.
To formalize these competitive neurons, we first define the competition class of a neuron $n_{i}$:
$\mathcal{C}(n_{i})=\bigl{\{}n_{i}^{k}\,\text{ where }\,\mathcal{B}_{r}(n_{i})=\mathcal{B}_{r}(n_{i}^{k})\bigr{\}},$ (15)
These neurons are mutually exclusive candidates to represent the context-dependent expression of a feature at a fixed spatial position. During each time step, after computing the pre-synaptic activities $\boldsymbol{a}_{j,t}^{\mathcal{I}}$ for $n_{j}\in\mathcal{C}(n_{i})$, we select a small winner set
$W_{t}(n_{i})=\mathrm{Comp}\bigl{(}\{\,\boldsymbol{a}_{j,t}^{\mathcal{I}}\mid n_{j}\in\mathcal{C}(n_{i})\}\bigr{)}.$ (16)
While various rules to determine the winning set can be used, we apply winner-takes-all selection in this work (for implementation details, see Section 3.2). Only neurons in $W_{t}(n_{i})$ are eligible to fire:
$y_{i,t}\;=\;\begin{cases}\mathcal{A}\bigl{(}\boldsymbol{a}_{i,t}^{\mathcal{I}},\mathcal{E}\bigr{)}&\text{if }n_{i}\in W_{t}(n_{i}),\\ 0&\text{otherwise.}\end{cases}$ (17)
By enforcing mutual exclusivity among neurons with overlapping receptive fields, competition ensures that different contexts activate distinct neuronal subsets.
@@ 3 Cooperative Network Architecture
Figure 2: Dynamics within the proposed CNA: The input image is fed into stage $S1$ to obtain feature activations. The feature activations are then processed by stage $S2$, together with activations from neurons within the same layer.
@@
The CNA comprises two processing stages, depicted as yellow squares in Figure 2. Stage 1 ($S1$) performs fixed (i.e., not learned) feature extraction directly from the image input. This fixed feature stage is used here for simplicity and to facilitate clearer diagnostics of the subsequent results by combining well-characterized, predefined features. However, these features could also be learned in future work, for example, through biologically inspired Hebbian learning schemes that have been shown to induce convolutional, Gabor-like filters (e.g., Miconi, 2021). Stage 2 ($S2$) implements the core principles of the described computational framework by learning net fragments.
The network dynamics unfold over discrete time steps $t=[0,\dots,T]$. Since the feature extraction stage ($S1$) employs four fixed filters, it yields a time-invariant activation pattern for a given (fixed) input. In contrast, the second stage ($S2$) exhibits dynamic temporal evolution, where internal representations are iteratively refined at each time step ($t\to t+1$). Specifically, we use a progressively increasing global attenuation signal that only allows neurons with a high pre-synaptic activity to remain active. The following sections discuss these two processing stages.
@@ 3.1 Stage 1: Feature Extraction
The input image is denoted as $\boldsymbol{x}\in\mathbb{R}^{C,H,W}$ (illustrated by the gray box in Figure 2), where $C$ indicates the number of image channels (111 for grayscale, 333 for RGB), while $H,W$ stand for the height and width of the image.
@@
To extract initial activation patterns, we use four fixed, non-trainable filters (details see Appendix A). These filters are implemented as standard convolutional kernels, used to obtain the binary activation pattern $\boldsymbol{y}^{(S1)}\in\{0,1\}^{C^{(S1)},H,W}$ as follows:
@@
$y_{i}=\Theta(\boldsymbol{a}_{i})=\begin{cases}1\quad\text{if }\boldsymbol{a}_{i}>0\\ 0\quad\text{otherwise}\end{cases}$ (19)
@@ 3.2 Stage 2: Net Fragments
$S2$ constructs net fragments from the binary feature activations $\boldsymbol{y}^{(S1)}$ of $S1$ and integrates them into coherent nets over time. To ensure stability, single neurons or small groups of neurons are suppressed through an attenuation mechanism, ensuring that only large-scale, consistent activation patterns persist.
@@
Each neuron $n^{(S2)}_{i}$ in stage 2 is connected in two ways: (1) Feedforward connections to the corresponding local neighborhood $\mathcal{B}_{r}(n^{(S1)}_{i})$ in $S1$, and (2) lateral connections to neighboring neurons within stage 2 itself, $\mathcal{B}_{r}(n^{(S2)}_{i})$, denoted as:
$\mathcal{B}_{r}\left(n^{(S2)}_{i}\right)$ $=\mathcal{B}_{r}\left(n^{(S1)}_{i}\right)\cup\left\{n_{j}^{(S2)}\mid d(n_{i}^{(S2)},n_{j}^{(S2)})\leq r\right\}$ (21)
$\mathcal{B}_{r}\left(n^{(S1)}_{i}\right)$ =$=\left\{n_{j}^{(S1)}\mid d(n_{i}^{(S2)},n_{j}^{(S1)})\leq r\right\}$
where $r$ defines the radius of lateral connections. The convolutional architecture realizes these neighborhood structures via convolutional filters, where the filter size directly determines the radius of both feedforward and lateral connections. Using convolutional filters not only implements the connectivity restrictions to local neighborhoods but also allows detection of the same pattern at various positions, enabling learning of net fragments independent of position.
@@
To determine the support a single neuron $n_{i,t}$ receives from the net fragment $\mathcal{F}_{t-1}(n_{i,t-1})$, we first calculate the pre-synaptic activity $\mathcal{I}$ by applying convolutional operations on the forward and lateral connections and then taking the sum of these operations:
@@
$\boldsymbol{a}^{\mathcal{I}(S2)}_{c_{\text{out}},t}$ $=\underbrace{\sum_{c_{\text{in}}=0}^{C^{(S1)}}\boldsymbol{W}^{(F)}_{c_{\text{out}},c_{\text{in}}}\star\boldsymbol{y}_{c_{\text{in}}}^{(S1)}}_{\textit{forward input }(\boldsymbol{a}^{\mathcal{I}(F)})}+\underbrace{\sum_{c_{\text{in}}=0}^{C^{(S2)}}\boldsymbol{W}^{(L)}_{c_{\text{out}},c_{\text{in}}}\star\boldsymbol{y}_{c_{\text{in}},t-1}^{(S2)}}_{\textit{lateral input }(\boldsymbol{a}^{\mathcal{I}(L)})}$ (22)
$\boldsymbol{a}^{\mathcal{I}(F)}_{c_{\text{out}}}$ $=\sum_{c_{\text{in}}=0}^{C^{(S1)}}\sum_{h=0}^{H-1}\sum_{w=0}^{W-1}\boldsymbol{W}^{(F)}_{c_{\text{out}},c_{\text{in}},h,w}\cdot\boldsymbol{y}^{(S1)}_{c_{\text{in}},h^{\prime}+h,w^{\prime}+w}$
$\boldsymbol{a}^{\mathcal{I}(L)}_{c_{\text{out}},t}$ $=\sum_{c_{\text{in}}=0}^{C^{(S2)}}\sum_{h=0}^{H-1}\sum_{w=0}^{W-1}\boldsymbol{W}^{(L)}_{c_{\text{out}},c_{\text{in}},h,w}\cdot\boldsymbol{y}^{(S2)}_{c_{\text{in}},h^{\prime}+h,w^{\prime}+w,t-1}$